What is the latitude of someone in Europe who sees the Sun at an altitude of 14.5° on December 22?
Answer #1

1 answer · Astronomy & Space

Best Answer

I presume you mean 'at noon' (meaning meridian passage, the time of maximum altitude).

The sun's declination is measured from the celestial equator.

The celestial equator is exactly 90 degrees from the celestial pole.

The celestial pole is exactly at the same altitude as the observer's latitude. Call this 'LAT'

Altitude of pole, over north horizon = LAT

Altitude of equator over north horizon = 90 + LAT (yes, you look up, past the zenith, and bend backwards).

This is the same as turning around (180 degrees) and looking over the south horizon: 180 - (90+LAT) = 90 - LAT

The celestial equator is (90 - LAT) over the south horizon.

The declination is measured from the celestial equator; when looking south, the declination is above the equator if it is north, and it is below the equator if it is south.

On December 22, the sun's declination is around 23.4 degrees south of the equator.

The altitude of the sun at noon on Dec. 22 is: (90 - LAT) - 23.4

And this is equal to 18.5 according to your question:

14.5 = 90 - LAT - 23.4

solve for + LAT (careful with the signs):

LAT = 90 - 14.5 -23.4 = 52.1

I hope it was helpful..